# Auto Arena of LLMs: Automating LLM Evaluations with Agent Peer-battles and Committee Discussions

### v3

@@ A.3 Inter-judge Agreement
As shown in Figure 4, the Cohen’s Kappa agreement [29] among judges before committee discussion is very low, averaging 0.10, which indicates slight agreement. We notice that weak model judges and strong model judges has an especially low agreement, such as GPT-4 and Llama -2. This shows that general model capabilities could result in significant performance gaps when used as judges.
A.4 Model selection for the main experiment
@@
GPT-4-0409-Turbo [34] Newest and Strongest in GPT-4 model family Proprietary
Claude-3-Haiku [2] Newest and Cheapest in Claude-3 model family Proprietary
Qwen1.5-72B-chat [3] Strongest in Qwen model family Qianwen LICENSE
Mixtral-8x7b-Instruct-v0.1 [23] Strongest in open-source Mistral small models Apache 2.0
@@
Yi-34B-Chat [1] Strongest in Yi Model Family Yi License
GPT-3.5-Turbo-0125 [32] Newest in GPT-3.5 Model Family Proprietary
Llama-2-70b-chat [38] Strongest in Llama Model Family Llama 2 Community
GPT-4o-2024-05-13 [33] Newly released model in GPT-4 Model Family Proprietary
Llama-3-70b-chat-hf [30] Newly released model in Llama Model Family Llama 3 Community
Reka-core-20240501 [37] Strongest model in Reka Model Family Proprietary
Command R Plus [15] Strongest model in Command R Model Family CC-BY-NC-4.0
mistral-large-2402 [16] Strongest model in mistral Family Proprietary
deepseek-llm-67b-chat [17] Strongest model in deepseek Family DeepSeek License
Wenxin-5 Famous Chinese model, not on Chatbot Arena Proprietary
GLM-4 Famous Chinese model, not on Chatbot Arena Proprietary
Minimax-abab6.5 Famous Chinese model, not on Chatbot Arena Proprietary
SenseChat-5 Famous Chinese model, not on Chatbot Arena Proprietary
@@
A.5 Comparison of baseline methods and Auto-Arena
@@
MMLU [22] Yes Static - Answer Accuracy
OpenLLM Leaderboard [6] Yes Static - Answer Accuracy
LC-AlpacaEval [19] Yes Static \$10 Single LLM (GPT-4)
MT-Bench [44] Yes Static \$10 Single LLM (GPT-4)
MT-Bench Hard [26] Yes Frequent Updates \$25 Single LLM (GPT-4)
Chatbot Arena [44] Yes Live Very High Humans
@@
Table 8 shows a comparison between benchmark evaluation methods and Auto-Arena. Compared to previous methods, the main advantage of Auto-Arena is the zero need for human dataset construction or intervention and the freshness of queries. Another main innovation compared to previous model-based systematic benchmarking procedures is using a committee of LLMs to discuss and vote for a final winner, which introduces diverse viewpoints.
For the evaluation cost, the costs of Auto-Arena are on the same scale as other benchmarks: We note that the primary experiment among 7 models costs around \$35 USD. Therefore, the estimated cost is \$5 per model. As models on the ranking board increase, the costs of conducting debates should grow slowly in log scale, which comes from conducting $nlog_{2}(n)$ pairings when adding 1 model to a ranking of $(n-1)$ models. The evaluation costs, however, shall remain the same as we use a committee of 5 LLMs at all times.

### v4

@@ Appendix F Inter-judge Agreement
As shown in Figure 11, the Cohen’s Kappa agreement (McHugh, 2012) among judges before committee discussion is very low, averaging 0.16, which indicates slight agreement. We notice that weak model judges and strong model judges has an especially low agreement, such as GPT-4 and Yi. This shows that general model capabilities could result in significant performance gaps when used as judges.
After the 1 round of communication, agreements significantly improved as the judges become convinced by more persuasive arguments. The average Cohen’s Kappa after discussion reaches 0.27, which indicates fair agreement.
Appendix G Model selection for the main experiment
@@
GPT-4-0409-Turbo (OpenAI et al., 2024) Newest and Strongest in GPT model family under GPT-4 Proprietary
GPT-4o-2024-05-13 (Openai, 2024b) Newly released model in GPT Model Family Proprietary
GPT-3.5-Turbo-0125 (Openai, 2024a) Newest ChatGPT version in the GPT Model Family Proprietary
Claude-3.5-Sonnet-20240620 (Anthropic, 2024) Newest in Claude model family under Claude-3.5 Proprietary
Claude-3-Haiku (Anthropic, 2024) Newest and Cheapest in Claude model family under Claude-3 Proprietary
Qwen/Qwen2-72B-Instruct (Bai et al., 2023) Representative of Qwen Model Family under Qwen-2 Proprietary
Qwen1.5-72B-chat (Bai et al., 2023) Representative of Qwen model family under Qwen-1.5 Qianwen LICENSE
Command R Plus (Cohere, 2024) Strongest model in Command R Model Family CC-BY-NC-4.0
Llama-3-70b-chat-hf (Meta, 2024) Representative of Llama Model Family under Llama-3 Llama 3 Community
Llama-2-70b-chat (Touvron et al., 2023) Representative of Llama Model Family under Llama-2 Llama 2 Community
Mixtral-8x7b-Instruct-v0.1 (Jiang et al., 2024) Strongest in open-source Mistral small models Apache 2.0
@@
Gemma-2-27b-it (Team et al., 2024a) Representative of the Gemma family Apache 2.0
Gemini-1.5-flash-exp-0827 (Team et al., 2024a) Cheapest in the Gemini-1.5 family Proprietary
Yi-34B-Chat (AI et al., 2024) Strongest in Yi Model Family on Chatbot Arena Yi License
Deepseek-LLM-67B-chat (DeepSeek-AI et al., 2024) Representative open-source model in Deepseek Family DeepSeek License
@@
Appendix H Comparison of baseline methods and Auto-Arena
@@
OpenLLM Leaderboard (Beeching et al., 2023) Yes Static - Answer Accuracy
MMLU (Hendrycks et al., 2021a) Yes Static - Answer Accuracy
GPQA (Rein et al., 2023) Yes Static - Answer Accuracy
LC-AlpacaEval (Dubois et al., 2024a) Yes Static \$10 Single LLM (GPT-4)
MT-Bench (Zheng et al., 2023) Yes Static \$10 Single LLM (GPT-4)
Arena Hard (Li* et al., 2024) Yes Frequent Updates \$25 Single LLM (GPT-4)
Chatbot Arena (Zheng et al., 2023) Yes Live Very High Humans
@@
Table 12 shows a comparison between benchmark evaluation methods and Auto-Arena. Compared to previous methods, the main advantage of Auto-Arena is the zero need for human dataset construction or intervention and the freshness of queries. Another innovation compared to previous model-based systematic benchmarking procedures is using a committee of LLMs to discuss and vote for a final winner, which introduces diverse viewpoints. The most important innovation of Auto-Arena is the peer-battle mechanism, which asks LLM agents to compete and debate with each other. The resulting evaluation on the multi-turn debate then becomes more in-depth, interactive, and comprehensive.
For the evaluation cost, the costs of Auto-Arena are on the same scale as other benchmarks: We note that the primary experiment among 9 models costs around \$45 USD. Therefore, the estimated cost is \$5 per model. As models on the ranking board increase, the costs of conducting debates should grow slowly in log scale, which comes from conducting $nlog_{2}(n)$ pairings when adding 1 model to a ranking of $(n-1)$ models. The evaluation costs, however, shall remain the same as we use a committee of 5 LLMs at all times.